# The Atacama Cosmology Telescope: DR6 Gravitational Lensing and SDSS BOSS cross-correlation measurement and constraints on gravity with the EG statistic

### v1

@@ V Blinding procedure
During development, we blinded the ACT DR6 CMB lensing map with an unknown blinding amplitude. We developed the null and consistency tests on the blinded map. We did not compare the measured curves to the theory during development. Furthermore, we finalized the estimator used in this work and the analysis choices like the binning scheme and cosmological ranges before unblinding.
After confirming the accuracy of the pipeline we performed a step-wise unblinding procedure. First, we used the amplitudes of the measurements to refine the galaxy bias values from their initial guesses of used during development. Then we recreated the galaxy simulations with the updated theory curves, reran the null and consistency tests on the unblinded maps, and confirmed they still show statistical consistency. The values for the null and consistency tests we show in the paper are the ones after unblinding. Next, we ran the pipeline and estimation on Planck and SDSS BOSS data. We published the developed estimator and results using Planck+SDSS BOSS in Wenzl et al. [26].
Finally, we unblined the ACT-based results and compared them to the GR predictions. After revealing the results we investigated the sensitivity of the results to various analysis choices and combined the two angular cross-power spectra to derive a combined constraint. We did not make significant changes to the analysis pipeline or the results after unblinding.
VI Angular cross-power spectrum measurement
VI.1 Analysis approach
In this section, we describe the analysis pipeline for the angular cross-power spectra measurements between ACT DR6 lensing and the SDSS LOWZ and CMASS samples including the angular cross-power spectrum estimation (Section VI.1.1), the creation of simulated maps (Section VI.1.2), the covariance estimation (Section VI.1.3), and modeling of observational and astrophysical effects: normalization correction in Section VI.1.4 and magnification bias in Section VI.1.5.
VI.1.1 Estimation of angular cross-power spectra
@@
We account for the partial completeness of each pixel in our maps by rescaling them and therefore use the default masked_on_input = False when creating NaMaster field objects. For the galaxy overdensity maps this rescaling is the inverse of the galaxy mask as described in Eq. 29. However, for maps, we use the inverse of the CMB mask squared which is an approximation of the mask of the lensing map. We test that this is a reasonable approximation for our purposes through our end-to-end pipeline test in Section VI.2.1.
@@
We use a log-spaced binning scheme. For computational accuracy, since different scales are correlated, the computations are performed over a sufficiently wide range limited by our map resolution: . We limit the cosmological analysis to throughout to be consistent with previous analyses [26, 71, 31]. Additionally, we limit the cosmological range to scales where the linear assumptions underlying the statistic are valid. This represents a comoving real space cutoff at around . For the effective redshift of the galaxy samples this results in an upper cutoff of for CMASS and for LOWZ where the exact values are chosen to be edges in our binning scheme. The data allows constraining of the angular cross-power spectrum to smaller scales beyond the linear regime. We perform our consistency tests in an extended validation range . The exact edges of the log-spaced binning scheme over the full extended range are given by where the bins are inclusive of the lower edge and exclusive of the upper edge.
Since this approach to get an unbiased estimate of the power spectrum assumes the underlying power spectrum to be a step function matching the bandpowers, we also need to correct the theory for the bandpower binning. This is done by applying the coupling matrix, then binning and then applying the inverse of the binned coupling matrix [70]. We apply this throughout to compare theory spectra to data and for calculating the magnification bias correction in Section VI.1.5.
VI.1.2 Simulated map realizations
The ACT DR6 lensing data products include 400 simulated CMB lensing map realizations [6]. These maps have undergone the same analysis pipeline as the data and therefore contain comparable error characteristics which makes them ideal for testing our analysis pipeline.
Using the input realizations that were used to create the simulated ACT maps we create 400 simulated Gaussian galaxy overdensity map realizations for both CMASS and LOWZ that have the expected correlation with the simulated CMB lensing maps, the expected auto-correlation, and the same noise level as the observed galaxies given by the shot noise estimate. As done for the data we create two versions of the galaxy maps for the auto- and cross-correlation where for the cross-correlation we account for the shifted from the additional reweighting we do for the cross-correlation only [Eq. 24]. This approach is described in detail in Wenzl et al. [26], building on previous work [72, 63].
Linear galaxy biases of for CMASS and for LOWZ are inferred by comparing the measured angular auto-power spectra of the galaxies to the amplitude of the input theory angular power spectra calculated for our fiducial cosmology. Before unblinding the data we used a preliminary value of [73, 74].
We use these simulated maps to estimate the baseline simulation-based covariance for the measured angular power spectra both for the cosmological results and for a range of null and constancy tests of the analysis. We also show, based on these simulated maps, that our analysis pipeline accurately recovers the input angular-power spectra (Section VI.2.1).
VI.1.3 Covariance estimation and consistency
@@
In addition to the covariance estimated from simulated realizations, we also estimate the covariance analytically. For this we use the angular power spectra and for our fiducial cosmology, the masks of the CMB lensing and galaxy maps, an estimate of the shot noise in the galaxy map and an estimate of the noise in the CMB lensing map. For the mask of the CMB lensing map, we use the square of the mask of the primary CMB observations which is an approximation we confirm to be sufficiently accurate through our input recovery test (Section VI.2.1). We use the method from García-García et al. [76] implemented in NaMaster as the NmtCovarianceWorkspace class to estimate the Gaussian covariances. This gives us analytic estimates of the covariance of as well as the covariance for and the cross-covariance between them.
@@
In Fig. 3 we compare the simulated covariance and analytic covariances showing the off-diagonals of the covariance as a correlation matrix. The correlation matrices are calculated from the covariances as . The off-diagonals of the covariance are consistent between the two methods within the statistical uncertainty of the simulation-based estimate. The most relevant correlations are between auto- and cross-power spectra at the same scale and for neighboring bins.
The amplitudes of the diagonals predicted by three different covariances approaches are also consistent as discussed in Section VI. As we will show in the comparison of the analytic and jackknife-based error estimates relative to the simulation-based estimates in the middle panels of Fig. 11, we find that both estimates are within expectations. This gives us confidence to use our simulation and analytic approaches as the covariance estimates.
VI.1.4 Normalization correction
@@
where and are the simulated input realizations of the lensing map, where the first one has the specific mask of the galaxy sample that we want to correlate with applied and the second one has the CMB mask squared applied. The restriction of the angular power spectrum to the overlap with the galaxy map considered ensures that we get the correction for the specific area relevant to the analysis. Using the square of the CMB mask an approximation of the mask for the derived lensing reconstruction. We confirm that this approximation is accurate for our purposes with our end-to-end pipeline test showing that we accurately recover the input spectra to the precision needed (Section VI.2.1). Note it is numerically advantageous to average over our set of sims before taking the ratio. Furthermore, the correlations are calculated only using Eq. 33, since we do not want to account for the coupling matrix in our normalization correction to avoid correcting for it twice. We implement this calculation using healpy.sphtfunc.anafast.
@@
We apply the normalization corrections to all simulated and observed reconstruction maps for our results and null tests. When comparing the North and South patches of SDSS we calculate the normalization correction separately for each. Since the normalization correction affects the amplitude of the cross-correlation and therefore , we evaluate to what degree our results and conclusion depend on the accuracy of the normalization correction in Section VII.3.
VI.1.5 Magnification bias
@@
VI.2 Consistency and null tests
We perform a range of null and consistency tests on our angular cross-power spectrum analysis pipeline to ensure its accuracy. We validate the pipeline by showing that we accurately recover the input based on simulations (Section VI.2.1). We further confirm our covariance estimates by considering the correlation of misaligned maps (Section VI.2.2). We perform a range of tests that are sensitive to foreground contamination to the result to build confidence in the cross-correlation analysis (Sections VI.2.3, VI.2.4 and VI.2.5). We show that the analysis is consistent between the SDSS North and South patches to justify combining them for the presented work (Section VI.2.6).
@@ VI.2 Consistency and null tests
where is the difference between the observed angular power spectra and the covariance for the difference. In some cases, the two measurements are independent, and therefore the cross-covariance can be neglected. However, in cases where the angular power spectra being compared are significantly correlated, it is crucial to account for the cross-covariance terms to not underestimate the value. Given the degrees of freedom (DOF) for the test, we calculate the probability of getting a value at least this large, referred to as probability to exceed (PTE). These PTE values are calculated as the integral of a distribution for the specific DOF from the measured value to infinity.
In the following paragraphs, we discuss the motivation and findings of individual null and consistency tests. The resulting PTE values are summarized in Table 1. In Section VI.2.7 we discuss the distribution of all PTE tests presented as part of this work.
@@
- MV, map level diff. | 6.2/11 | 0.86 | 12.6/16 | 0.70 | 16.1/8 | 0.04 | 23.2/16 | 0.11 | |||||||||
- TT, map level diff. | 13.4/11 | 0.27 | 17.5/16 | 0.35 | 12.7/8 | 0.12 | 20.4/16 | 0.20 | |||||||||
- MV, band power diff. | 15.0/11 | 0.18 | 20.3/16 | 0.21 | 2.7/8 | 0.95 | 9.3/16 | 0.90 | |||||||||
- TT, band power diff. | 9.9/11 | 0.54 | 12.5/16 | 0.71 | 6.8/8 | 0.56 | 10.3/16 | 0.85 | |||||||||
@@
North vs South | Combined analysis | 8.6/11 | 0.66 | 15.3/16 | 0.50 | 2.4/8 | 0.97 | 7.5/16 | 0.96 |
VI.2.1 Accurate input recovery
We test if the analysis pipeline accurately recovers the input angular cross-power spectrum for our fiducial cosmology. We use the 400 simulated ACT CMB lensing and galaxy map realizations discussed in Section VI.1.2 and apply our analysis pipeline to each combination. This test covers our full analysis pipeline including the normalization correction (see Section VI.1.4) and also tests if the approximation of using the CMB mask squared as an approximation for the mask of the CMB lensing map is sufficiently accurate for the analysis.
In Fig. 4 we show the mean estimated angular cross-power spectra from all simulations compared to the input. We test the difference between the mean recovered measurement and the input for statistical consistency. Here the input does not have an uncertainty, therefore the covariance in Eq. 43 is given by scaling the covariance of an individual measurement . In the cosmological analysis ranges we find PTE values of and for ACT CMASS and ACT LOWZ and using the baseline simulation-based covariance. This shows that within statistical expectations we recover the input signal. We can also translate this into the absolute deviation from the input for our pipeline test. This depends on the number of simulations used. We fit the theory curve to the mean recovered signal with a free amplitude , with perfect recovery yielding . We find for values of and for ACT CMASS and ACT LOWZ respectively. We conservatively account for the level at which we validated the amplitude in our systematic error budget. We use 95% upper limits of and respectively which are small compared to the statistical uncertainties of the estimates.
@@ VI.2.1 Accurate input recovery
VI.2.2 Misaligned correlation null test
@@
VI.2.3 Websky foreground bias estimate
@@
We test the foreground mitigation scheme on the Websky simulation [64, 65]. Specifically, we run the ACT reconstruction pipeline on the foreground-only CMB temperature map that models the contributions of CIB, kSZ, tSZ, and radio point sources to the observed CMB temperature anisotropy. All of these act as contaminants in the reconstruction of the CMB lensing signal. By applying the reconstruction pipeline on the foreground-only map we find the residual signal these foregrounds leave in the reconstructed map. If we find these foreground residuals to be correlated with a galaxy map constructed from the same simulation this would indicate that the foregrounds would contaminate our observed cross-correlation measurement. When instead we find that the correlation is small compared to our measurement uncertainty it builds confidence that the reconstruction approach sufficiently mitigates foreground contamination from extragalactic sources.
The Websky simulation includes a halo catalog that we can leverage to build galaxy samples that match the properties of our galaxy samples to test the mitigation for a comparable sample. We construct a CMASS-like galaxy catalog from the Websky halo catalog by applying an HOD which is tuned to closely match the CMASS galaxy sample [80, 81], following previous approaches [e.g. 82, 31]. The HOD has the form
with best fit values , , , and [82]. For each halo, we sample observed galaxies based on and place them in the center of the halo. For each halo that has a central galaxy, we Poisson sample satellite galaxies based on and placed in a random direction from the center with the radius sampled from a Navarro et al. [83] (NFW) profile using halo concentration parameters from Duffy et al. [84]. Finally, we subsample the galaxies to match the number density and redshift distribution of the CMASS sample. This results in a CMASS-like galaxy catalog which is correctly correlated with the other data products of the Websky simulation. For LOWZ we follow a similar approach implementing the HOD from Parejko et al. [Table 3, column Mean Full in 74] with a slightly different functional form but otherwise proceed the same way to create a galaxy catalog. We also note that for the redshift range 0 to 1 the minimum halo mass captured by Websky is sufficient for our purposes for both samples.
@@
VI.2.4 90-150 GHz null test
To test for contamination of our estimate from foregrounds we can use tests on the data itself in addition to the simulation-based test discussed in Section VI.2.3. One way is to purposely null the cosmological signal in the CMB lensing map and test if there is any residual correlation with the galaxy maps which would indicate some foreground contamination.
By taking the difference between the and maps we null the CMB lensing information in the primary CMB. This also nulls frequency-independent components like kSZ but other foreground effects can still be in the map. By running the lensing reconstruction pipeline on the difference we create a null map that should be uncorrelated with the galaxy maps which we can then test. We can also run the reconstruction on the individual frequency maps and then take the difference in the final bandpowers. Finally, we can limit this test to the temperature data alone and check if there is a difference to the baseline minimum variance combination of temperature and polarization information.
@@
The PTE value for the 90–150 GHz null test at the map level using both temperature and polarization and for ACT LOWZ is marginally below 0.05 with a value of PTE=0.04. Given the number of null tests performed in this analysis having a PTE value this low is not surprising but it motivates a check for those results to identify if there is a concern. The same test for only the temperature information passes and the test at the bandpower level also passes further indicating this could be a statistical outlier. We find this null test does not fail if we add or remove a bin at the small-scale end. When adding a bin at the large-scale end the test still marginally fails, when removing one it passes. The test does pass for our full extended validation range. This dependence on the test range could be an indication that this is indeed a chance failure. In the plot (orange points in the top right panel of Fig. 7), we do not see clear outliers and there is no scale-dependent trend visible. Overall we conclude that this PTE value of 0.04 is within statistical expectation and we did not identify a reason for concern. In Section VI.2.7 we summarize all PTE value tests in this analysis and evaluate whether there are more low PTE values as expected for the number of tests performed.
Overall this demonstrates the robustness of the analysis, and indicates that the difference map does not have any significant residuals correlated with the galaxy maps.
VI.2.5 Consistency with CIB-deprojection approach

### v2

@@ IV Blinding procedure
During development, we blinded the ACT DR6 CMB lensing map with an unknown blinding amplitude (see Qu et al. [21] for additional details on the blinding approach). We developed the null and consistency tests on the blinded map. We did not compare the measured curves to the theory during development. Furthermore, we finalized the estimator used in this work and the analysis choices like the binning scheme and cosmological ranges before unblinding.
After confirming the accuracy of the pipeline we performed a step-wise unblinding procedure. First, we used the amplitudes of the measurements to refine the galaxy bias values from their initial guesses of used during development. Then we recreated the galaxy simulations with the updated theory curves, reran the null and consistency tests on the unblinded maps, and confirmed they still show statistical consistency. The values for the null and consistency tests we show in the paper are the ones after unblinding. We use an estimator developed on SDSS and BOSS data from Wenzl et al. [27].
Finally, we unblined the ACT-based results and compared them to the GR predictions. After revealing the results we investigated the sensitivity of the results to various analysis choices and combined the two angular cross-power spectra to derive a combined constraint.
V Angular cross-power spectrum measurement
V.1 Analysis approach
V.1.1 Estimation of angular cross-power spectra
@@
We account for the partial completeness of each pixel in our maps by rescaling them and therefore use the default masked_on_input = False when creating NaMaster field objects. For the galaxy overdensity maps this rescaling is the inverse of the galaxy mask as described in Eq. 23. However, for maps, we use the inverse of the CMB mask squared which is an approximation of the mask of the lensing map. We test that this is a reasonable approximation for our purposes through our end-to-end pipeline test in Section V.2.1.
@@
We use a log-spaced binning scheme. For computational accuracy, since different scales are correlated, the computations are performed over a sufficiently wide range limited by our map resolution: . We limit the cosmological analysis to throughout as ACT is not sensitive to the largest scales. The cutoff value is chosen to be consistent with previous analyses [27, 72, 32]. Additionally, we limit the cosmological range to scales where the linear assumptions underlying the statistic are valid. This represents a comoving real space cutoff at around . For the effective redshift of the galaxy samples this results in an upper cutoff of for CMASS and for LOWZ where the exact values are chosen to be edges in our binning scheme. The data allows constraining of the angular cross-power spectrum to smaller scales beyond the linear regime. We perform our consistency tests in an extended validation range . The exact edges of the log-spaced binning scheme over the full extended range are given by [40, 48, 59, 71, 87, 106, 129, 157, 191, 233, 283, 345, 420, 511, 622, 757, 922] where the bins are inclusive of the lower edge and exclusive of the upper edge.
Since this approach to get an unbiased estimate of the power spectrum assumes the underlying power spectrum to be a step function matching the bandpowers, we also need to correct the theory for the bandpower binning. This is done by applying the coupling matrix, then binning and then applying the inverse of the binned coupling matrix [71]. We apply this throughout to compare theory spectra to data and for calculating the magnification bias correction in Section V.1.4.
V.1.2 Covariance estimation and consistency
@@
In addition to the covariance estimated from simulated realizations, we also estimate the covariance analytically. For this we use the angular power spectra and for our fiducial cosmology, the masks of the CMB lensing and galaxy maps, an estimate of the shot noise in the galaxy map and an estimate of the noise in the CMB lensing map. For the mask of the CMB lensing map, we use the square of the mask of the primary CMB observations which is an approximation we confirm to be sufficiently accurate through our input recovery test (Section V.2.1). We use the method from García-García et al. [74] implemented in NaMaster as the NmtCovarianceWorkspace class to estimate the Gaussian covariances. This gives us analytic estimates of the covariance of as well as the covariance for and the cross-covariance between them.
@@
We compare the simulated covariance and analytic covariances showing the off-diagonals of the covariance as a correlation matrix. The correlation matrices are calculated from the covariances as . The off-diagonals of the covariance are consistent between the two methods within the statistical uncertainty of the simulation-based estimate. The most relevant correlations are between auto- and cross-power spectra at the same scale and for neighboring bins.
The amplitudes of the diagonals predicted by three different covariances approaches are also consistent as discussed in Section V. As we will show in the comparison of the analytic and jackknife-based error estimates relative to the simulation-based estimates in the middle panels of Fig. 8, we find that both estimates are within expectations. This gives us confidence to use our simulation and analytic approaches as the covariance estimates.
V.1.3 Normalization correction
@@
where and are the simulated input realizations of the lensing map, where the first one has the specific mask of the galaxy sample that we want to correlate with applied and the second one has the CMB mask squared applied. The restriction of the angular power spectrum to the overlap with the galaxy map considered ensures that we get the correction for the specific area relevant to the analysis. Using the square of the CMB mask an approximation of the mask for the derived lensing reconstruction. We confirm that this approximation is accurate for our purposes with our end-to-end pipeline test showing that we accurately recover the input spectra to the precision needed (Section V.2.1). Note it is numerically advantageous to average over our set of sims before taking the ratio. Furthermore, the correlations are calculated only using Eq. 27, since we do not want to account for the coupling matrix in our normalization correction to avoid correcting for it twice. We implement this calculation using healpy.sphtfunc.anafast.
@@
We apply the normalization corrections to all simulated and observed reconstruction maps for our results and null tests. When comparing the North and South patches of SDSS we calculate the normalization correction separately for each. Since the normalization correction affects the amplitude of the cross-correlation and therefore , we evaluate to what degree our results and conclusion depend on the accuracy of the normalization correction in Section VI.3.
V.1.4 Magnification bias
@@
V.2 Consistency and null tests
We perform a range of null and consistency tests on our angular cross-power spectrum analysis pipeline to ensure its accuracy. We validate the pipeline by showing that we accurately recover the input based on simulations (Section V.2.1). We further confirm our covariance estimates by considering the correlation of misaligned maps (Section V.2.2). We perform a range of tests that are sensitive to foreground contamination to the result to build confidence in the cross-correlation analysis (Sections V.2.3, V.2.4 and V.2.5). We show that the analysis is consistent between the SDSS North and South patches to justify combining them for the presented work (Section V.2.6).
@@ V.2 Consistency and null tests
where is the difference between the observed angular power spectra and the covariance for the difference. In some cases, the two measurements are independent, and therefore the cross-covariance can be neglected. However, in cases where the angular power spectra being compared are significantly correlated, it is crucial to account for the cross-covariance terms to not underestimate the value. Given the degrees of freedom (DOF) for the test, we calculate the probability of getting a value at least this large, referred to as probability to exceed (PTE). These PTE values are calculated as the integral of a distribution for the specific DOF from the measured value to infinity. Note that for Gaussian uncertainty a PTE value of 0.68 corresponds approximately to a difference and a PTE value of 0.05 approximately to . For previous applications of PTE values see for example Qu et al. [21], Farren et al. [64], Wenzl et al. [27].
In the following paragraphs, we discuss the motivation and findings of individual null and consistency tests. The resulting PTE values are summarized in Table 1. In Section V.2.7 we discuss the distribution of all PTE tests presented as part of this work.
@@
| - MV, map level diff. | 6.2/11 | 0.86 | 12.6/16 | 0.70 | 16.1/8 | 0.04 | 23.2/16 | 0.11 | |||||||||
| - TT, map level diff. | 13.4/11 | 0.27 | 17.5/16 | 0.35 | 12.7/8 | 0.12 | 20.4/16 | 0.20 | |||||||||
| - MV, band power diff. | 15.0/11 | 0.18 | 20.3/16 | 0.21 | 2.7/8 | 0.95 | 9.3/16 | 0.90 | |||||||||
| - TT, band power diff. | 9.9/11 | 0.54 | 12.5/16 | 0.71 | 6.8/8 | 0.56 | 10.3/16 | 0.85 | |||||||||
@@
| North vs South | Combined analysis | 8.6/11 | 0.66 | 15.3/16 | 0.50 | 2.4/8 | 0.97 | 7.5/16 | 0.96 | ||||||||
V.2.1 Accurate input recovery
We test if the analysis pipeline accurately recovers the input angular cross-power spectrum for our fiducial cosmology. We use the 400 simulated ACT CMB lensing and galaxy map realizations discussed in Section III.3 and apply our analysis pipeline to each combination. This test covers our full analysis pipeline including the normalization correction (see Section V.1.3) and also tests if the approximation of using the CMB mask squared as an approximation for the mask of the CMB lensing map is sufficiently accurate for the analysis.
We have tested that the mean angular cross-power spectra from all simulations recover the input signal. Here the input does not have an uncertainty, therefore the covariance in Eq. 37 is given by scaling the covariance of an individual measurement . In the cosmological analysis ranges we find PTE values of and for ACT CMASS and ACT LOWZ and using the baseline simulation-based covariance. This shows that within statistical expectations we recover the input signal. We can also translate this into the absolute deviation from the input for our pipeline test. This depends on the number of simulations used. We fit the theory curve to the mean recovered signal with a free amplitude , with perfect recovery yielding . We find for values of and for ACT CMASS and ACT LOWZ respectively. We conservatively account for the level at which we validated the amplitude in our systematic error budget. We use 95% upper limits of and respectively which are small compared to the statistical uncertainties of the estimates.
@@ V.2.1 Accurate input recovery
V.2.2 Misaligned correlation null test
@@
V.2.3 Websky foreground bias estimate
@@
We test the foreground mitigation scheme on the Websky simulation [62, 63]. Specifically, we run the ACT reconstruction pipeline on the foreground-only CMB temperature map that models the contributions of CIB, kSZ, tSZ, and radio point sources to the observed CMB temperature anisotropy. All of these act as contaminants in the reconstruction of the CMB lensing signal. Applying the reconstruction pipeline on the foreground-only map gives the residual signal these foregrounds leave in the reconstructed map. If these foreground residuals are correlated with a galaxy map constructed from the same simulation this would indicate that the foregrounds would contaminate our observed cross-correlation measurement. Therefore a null result builds confidence that the reconstruction approach sufficiently mitigates foreground contamination from extragalactic sources.
The Websky simulation includes a halo catalog that we can leverage to build galaxy samples that match the properties of our galaxy samples to test the mitigation for a comparable sample. We construct a CMASS-like galaxy catalog from the Websky halo catalog by applying an HOD which is tuned to closely match the CMASS galaxy sample [78, 79], following previous approaches [e.g. 80, 32]. The HOD has the form
with best fit values , , , and [80]. For each halo, we sample observed galaxies based on and place them in the center of the halo. For each halo that has a central galaxy, we Poisson sample satellite galaxies based on and placed in a random direction from the center with the radius sampled from a Navarro et al. [81] (NFW) profile using halo concentration parameters from Duffy et al. [82]. Finally, we subsample the galaxies to match the number density and redshift distribution of the CMASS sample. This results in a CMASS-like galaxy catalog which is correctly correlated with the other data products of the Websky simulation. For LOWZ we follow a similar approach implementing the HOD from Parejko et al. [Table 3, column Mean Full in 68] with a slightly different functional form but otherwise proceed the same way to create a galaxy catalog. We also note that for the redshift range 0 to 1 the minimum halo mass captured by Websky is sufficient for our purposes for both samples.
@@
V.2.4 90-150 GHz null test
To test for contamination of our estimate from foregrounds we can use tests on the data itself in addition to the simulation-based test discussed in Section V.2.3. One way is to purposely null the cosmological signal in the CMB lensing map and test if there is any residual correlation with the galaxy maps which would indicate some foreground contamination.
By taking the difference between the and maps we null the CMB lensing information in the primary CMB. This also nulls frequency-independent components like kSZ but other foreground effects can still be in the map. By running the lensing reconstruction pipeline on the difference we create a null map that should be uncorrelated with the galaxy maps which we can then test. We can also run the reconstruction on the individual frequency maps and then take the difference in the final bandpowers. Finally, we can limit this test to the temperature data alone and check if there is a difference compared to the baseline minimum variance combination of temperature and polarization information.
@@
The PTE value for the 90–150 GHz null test at the map level using both temperature and polarization and for ACT LOWZ is marginally below 0.05 with a value of PTE=0.04. Given the number of null tests performed in this analysis having a PTE value this low is not surprising but it motivates a check for those results to identify if there is a concern. The same test for only the temperature information passes and the test at the bandpower level also passes further indicating this could be a statistical outlier. We find this null test does not fail if we add or remove a bin at the small-scale end. When adding a bin at the large-scale end the test still marginally fails, when removing one it passes. The test does pass for our full extended validation range. This dependence on the test range could be an indication that this is indeed a chance failure. In the plot (orange points in the top right panel of Fig. 5), we do not see clear outliers and there is no scale-dependent trend visible. Overall we conclude that this PTE value of 0.04 is within statistical expectation and we did not identify a reason for concern. In Section V.2.7 we summarize all PTE value tests in this analysis and evaluate whether there are more low PTE values as expected for the number of tests performed. The difference map does not show a statistically significant correlation with the galaxy maps.
V.2.5 Consistency with CIB-deprojection approach